OpenStudy (geekfromthefutur):

Choose the equation that represents the graph below. Graph of a line passing through points 0 comma 4 and 3 comma 0 y = -negative three fourthsx + 4 y = -negative four thirdsx + 4 y = three fourthsx - 4 y = 4 thirdsx - 4

OpenStudy (anonymous):

Sorry not exactly good at this....

OpenStudy (geekfromthefutur):

heres the graph

OpenStudy (geekfromthefutur):

its ok im not to smart with math its difficult for me lol

OpenStudy (anonymous):

A maybe???

OpenStudy (anonymous):

y = -negative four thirdsx + 4

geerky42 (geerky42):

slope intercept form of equation looks like this: y = mx+b first we need to find slope of this line. do you know how to?

OpenStudy (geekfromthefutur):

thx matricked

OpenStudy (geekfromthefutur):

hey @matricked can u help me with another question?

OpenStudy (geekfromthefutur):

A student needs to make a rectangular cardboard piece that is 12 inches long. The area of the cardboard piece should be between 60 square inches to 300 square inches. The function f(w) relates the area of the cardboard, in square inches, to the width w in inches. Which of the following shows a reasonable domain for f(w)? 2 ≤ w ≤ 5 5 ≤ w ≤ 25 25 ≤ w ≤ 60 60 ≤ w ≤ 300

OpenStudy (anonymous):

if possible..

OpenStudy (anonymous):

5 ≤ w ≤ 25

OpenStudy (anonymous):

12*5=60 and 12*25=300

OpenStudy (geekfromthefutur):

oh ok thank you bro (:

OpenStudy (anonymous):

yw